OpenStudy (anonymous):

4x+6y=14 -2x+2y=-2 (put in (x,y) form)

OpenStudy (btaylor):

For the first one: 6y=14-4x y=14/6-4x/6 y=-2x/3+7/3 OR y=(-2x+7)/3 For the second one: 2y=2x-2 y=x-1